OpenStudy (anonymous):

13/32 plus 5/8 plus 1/2

OpenStudy (anonymous):

all have a common denominator of 32

OpenStudy (anonymous):

49/32

OpenStudy (anonymous):

\[\frac{13}{32}+\frac{20}{32}+\frac{16}{32}\]

OpenStudy (anonymous):

\[=\frac{49}{32}\]